OpenStudy (rayford):

openstudy what is the simplified form of y^2+7y+12/y^2-3y-18

OpenStudy (tkhunny):

Do you mean \[\frac{y^{2}+7y+12}{y^{2}-3y-18}\] That is NOT what you have written. Remember your Order of Operations. Factor Numerator and Denominator.

OpenStudy (rayford):

yes... i'm sorry.

OpenStudy (rayford):

may you please solve this so i can review it? i am very strong at reviewing.

OpenStudy (tkhunny):

That's funny. You need to know how to factor. Please demonstrate.

OpenStudy (rayford):

what do you mean? i don't know how :/ that's why i'm asking

OpenStudy (raden):

well, y^2+7y+12 = (y+3)(y+4) and y^2-3y-18 = (y-6)(y+3) now, just simplify it

OpenStudy (rayford):

u just made me feel really stupid, i don't know how to simplify..however, i just understood what you just said. what would be the simplified answer?

OpenStudy (raden):